ok if m=6/7,(5,-1) and we look at it with the equation y--y1=m(x-x1) so it would be y-(-1)=(6/7)(x-5) that then goes on to y+1=6/7x-30/7 i understand all of this just fine what confuses me is on the last step it says i must subtract -1 from both sides so how does the answer come out to be 37/7? final equation shows y=6/7x- 37/7 i dont get how subtracting -1 from both sides produces 37/7 from 30/7 please explain

Answer by rfer(16322)

y=mx+b
-1=6/7*5+b
-1=30/7+b
-1-30/7=b
b=-37/7
--------
y=6/7x-37/7 this is the answer
-------------
-1 becomes -7/7 and is added to -30/7 to get the -37/7, you add them because they are both minus numbers.